1. Calculate the binding energy (kJ/mol) for 84Kr.

1H = 1.00785 g/mol, 1n = 1.008665 g/mol, 84Kr = 83.91151 g/mol
Answer #1

protons is Kr = 36

neutrons = 84-36 = 48

The Kr atom is composed of 36 protons, 48 neutrons and 36 electrons.
The mass of 36 protons = 36 x 1.007277 amu
The mass of 48 neutrons = 48 x 1.008665 amu

The sum of the masses is 84.677892 amu
The actual mass, as measured by a mass spectrometer, is 83.911507 amu.
The difference ( 84.677892 - 83.911507 ) is 0.766385 amu

Convert the mass defect into kg.

(0.766385 amu/nucleus)(1.6606 x 10-27 kg/amu) = 1.2721991 x 10-27 kg/nucleus

Convert this mass into energy using delta E = delta m c2, where c = 2.9979 x 108 m/s

E = 1.1433 x 10-10 J/nucleus

To convert the energy to kJ/mol of nuclei

(1.1433 x 10-10 J/nucleus)(1 kJ/1000 J)(6.022 x 1023 nuclei/mol) = 6.9 x 1010 kJ/mol of nuclei